charge Q are uniformly distibuted ontwo concentric hollow sphare are radious r and R where(R>r). and surface charge density is same. find potential on commmon centre.

Grade:12

1 Answers

Arun
6 years ago
charge density=q/4(3.14)r=Q/4(3.14)R.
ie, q=Q*r*r/R*R.
potential at center of shell=kq/r+kQ/R
kQ/R(r/R+1)
Q=charge density*4(3.14)R*R
on simpilfying,
V=charge density*(r+R)/epsilon